OpenStudy (anonymous):

A gym offers regular memberships for $80 per month and off-peak memberships for $60 per month. Last month, the gym sold a total of 420 memberships for a total of $31,100. How many memberships sold were regular memberships?

OpenStudy (anonymous):

A) 125 B) 140 C) 235 D) 295

OpenStudy (anonymous):

Anyone who helps gets a medal (I think its A- 125 But I just want to be sure) @Kainui @undeadknight26 @campbell_st @Firejay5 @Nnesha

undeadknight26 (undeadknight26):

Its not 125 sorry bro ._.

OpenStudy (anonymous):

You sure??? What is it? =/ @undeadknight26

OpenStudy (anonymous):

Think you could do a little more numbers and a little less letters? xD?... I square Math was easy until thy added letters =/ @campbell_st

undeadknight26 (undeadknight26):

I wont tell you but: 125 x 80 = 10000 429 - 125 = 295 295 x 60 = 17700 17700 + 10000 = 27700

OpenStudy (anonymous):

you got 125 is membership for off peak. Now you have to substract it from 420 to get regular membership

undeadknight26 (undeadknight26):

This is pretty much what @campbell_st told you but with letters.

OpenStudy (anonymous):

Uggg 295??? @campbell_st Im in a hurry man

OpenStudy (anonymous):

Can you give me a problem like \[296-1\approx29\] =D

OpenStudy (anonymous):

That's embarrassing 295* e.e

OpenStudy (anonymous):

Guys look I the middle if this huge test. Look I know your not supposed to d this but if you could just give me the answer I would appreciate it. I haven't slept in two days Im just so tired...

OpenStudy (anonymous):

@Nnesha Please Im so tired ;-;

Nnesha (nnesha):

go to sleep :-)

Nnesha (nnesha):

then tomorrow start working on this now u not gonna understand anything take some rest :)

OpenStudy (anonymous):

You don't understand I cant

OpenStudy (anonymous):

I have to finish this tonight D=

OpenStudy (anonymous):

@shrutipande9 SOS

Nnesha (nnesha):

then do what campbell_stis saying \(\Huge\ddot{\Large\frown}\) not trying to be mean

OpenStudy (anonymous):

I would if I knew how lol

OpenStudy (anonymous):

@paki plz bb

OpenStudy (anonymous):

WAIT @Nnesha YOU CAN DO THIS YOUR A HUMAN CALCULATOR :O

OpenStudy (anonymous):

I wish I was a human calculator :3

Nnesha (nnesha):

let me remind u something this is open study not a calculator where you just have to click on post button and then you get answer think about that person who is trying to help you please try to understand put a little bit effort and do that he is saying :)

OpenStudy (anonymous):

Yes well with all do respect this one exam holds the key to the cllegde I get into and I haven't slept in over 2 nights because of it, so that's all im sayin is maybe you could say 3 numbers for me and help someones life :)

OpenStudy (anonymous):

\[\prime =m \alpha \sqrt{\sum_{\approx}^{?}}m \zeta2\]

OpenStudy (anonymous):

@TheGh0st @sarahprincess98 @jtryon @mathmath333 @ganeshie8 @eric_d

OpenStudy (anonymous):

ill take the general directions

OpenStudy (anonymous):

@EinsteinPart2 if someone is already helping you, you don't need to tag everyone just fyi :)

OpenStudy (anonymous):

No one was helping me -.- @jtryon

OpenStudy (anonymous):

@campbell_st was helping you, he is not going to give you the answer but at least he was pointing you in how to solve the problem

OpenStudy (anonymous):

Is the answer dun dun dun,,,, 295 :O!??!?!!?

OpenStudy (anonymous):

Heller

OpenStudy (anonymous):

Hello hello hello?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

I will help you. Which part don't you get?

OpenStudy (anonymous):

I GOTIT WRONG '

OpenStudy (anonymous):

name your variable. r= is number of regular memebership 0= is off peak of membership

OpenStudy (anonymous):

now total number membership is 420 r + o = 420 You got it

OpenStudy (anonymous):

Ok you see this is the time where you just give it up and give me the answer :L

OpenStudy (anonymous):

WAIT SO IT IS 250

OpenStudy (anonymous):

now total amount is $ 31,100, if you multiply number of membership of each kind to its unit cost you will get the equation 80r + 60 O = 31,100

OpenStudy (anonymous):

I have almost been awake for 3 days straight. I am begging you. Please have mercy on me and tell me if I was right, just say yes or no

OpenStudy (anonymous):

no. the answer is 295

OpenStudy (anonymous):

you know how to get it? let me know i can explain

OpenStudy (anonymous):

Thank you s much I love you

Nnesha (nnesha):

:/

OpenStudy (anonymous):

by solving the two equation above you get o = 125, then solve for r , r = 420 - 125 = 295

OpenStudy (anonymous):

Nether if you suspend anyone suspend me

OpenStudy (anonymous):

You are welcome. I am sorry for all the frustration. They will kick me out

OpenStudy (anonymous):

Any more question?

OpenStudy (anonymous):

Nesha* If you kick anyone out kick me out he was just trying to help

OpenStudy (anonymous):

i don't understand why they try to kick me out. I try to explain in the way that teacher at school will do the same

OpenStudy (anonymous):

And no im good thank you. I LOVE YOU GUYS GOONIGH BEDTIME BED OMFG YS BED SLEEP YES BED DED YES YES BED SLEEP YES YES SLEEP YES BED SLEEP YES BED YES SLEEP BED GOODNIGHT I LOVE YOU GUYS

Nnesha (nnesha):

there is a difference between just explain question and explain with an a answer that's a problem do you thing that he did understand that :/

OpenStudy (anonymous):

hey you can kick me out. I try to explain in the way that I know how. Don't torture people, he also try his best to get the answer, but it was wrong.

OpenStudy (anonymous):

I understand it. You have to use two different variables and the then plug them in for the substituted number and the substitutive value then after doing the math and analyzing the data you reach your conclusion which in this case was 295. If it was for Phuong I wouldn't have known that, therefore you have no right to kick either of us out. Night babies <3

OpenStudy (anonymous):

btw, you can kick me out. I don't need help, I only try to help people

Nnesha (nnesha):

i'm not trying to be rude 2 users already tried to help him but the problem was that they want him to do some work so his job was to put some effort instead of tagging other users

OpenStudy (anonymous):

Thanks Einstein. You understood. I am glad.

OpenStudy (anonymous):

He did try. Don't you see it. He got the answer 125, but he missed the next step. So, I help him. Feel free to kick me out. I don't care. I only want to help student

OpenStudy (anonymous):

I UNDERSTOOD if you take this ANY further you are abusing your job. Period. He did his job and I know fully understand the concept. The End ;) Night nigt

OpenStudy (anonymous):

obviously he understood what I try to say, but he did not understand other two users

Nnesha (nnesha):

yeah i saw that ... ;)

OpenStudy (firejay5):

@EinsteinPart2 Do you need any extra help on this problem or do you think you got it now?